Astronauts in the International Space Station must work out every day to counteract the effects of weightlessness. Researchers have investigated if riding a stationary bicycle while experiencing artificial gravity from a rotating platform gives any additional cardiovascular benefit. What frequency of rotation, in rpm, is required to give an acceleration of 1.4g to an astronaut's feet, if her feet are 1.1m from the platform's rotational axis?
Answer #1

The linear acceleration of a body in rotation is given by

a=\omega^2r

Over here, the rotating body is the astronaut's foot. Let us take g=9.81m/s

Substituting the given values, we get

1.4 x 9.81m/s= wx 1.1

Solving for angular velocity \omega , we get

w = 3.53rad/s

To express this in RPM, we multiply by 60 ( sixty seconds in a minute ), and divide by 2\pi ( 2\pi radians in full rotation ).

Therefore

The frequency of rotation in RPM

w x 60 -= 33.71RPM 27